write an equation with the given points parallel to the given line
(6,8) x+6y=5

Answer by toidayma(44)

The line paralleled to the line x + 6y = 5 has the equation of x + 6y + c = 0
Since this line pass the point (6,8), therefore: 6 + 6*8 + c = 0 or c = -54.
So, the finding line's equation is: x + 6y - 54 = 0.